if x^4 = x^2 - 72 find x
Answer by fractalier(6550)

We can transform
x^4 = x^2 - 72
into
x^4 - x^2 + 72 = 0
Now let
y = x^2
and we get
y^2 - y + 72 = 0
which is not factorable.
Are you sure this correct?
You sure the problem didn't say +72 above?
If this is indeed correct, apply the quadratic formula and then take the square root of each of your roots for y, yielding four solutions.